# Customizing Spider Silk: Generative Models with Mechanical Property Conditioning for Protein Engineering

### v1

@@ Abstract
The remarkable mechanical properties of spider silk, including its tensile strength and extensibility, are primarily governed by the repetitive regions of the proteins that constitute the fiber, the major ampullate spidroins (MaSps). However, establishing correlations between mechanical characteristics and repeat sequences is challenging due to the intricate sequence-structure-function relationships of MaSps and the limited availability of annotated datasets. In this study, we present a novel computational framework for designing MaSp repeat sequences with customizable mechanical properties. To achieve this, we developed a lightweight GPT-based generative model by distilling the pre-trained ProtGPT2 protein language model. The distilled model was subjected to multilevel fine-tuning using curated subsets of the Spider Silkome dataset. Specifically, we adapt the model for MaSp repeat generation using 6,000 MaSp repeat sequences and further refine it with 572 repeats associated with experimentally determined fiber-level mechanical properties. Our model generates biologically plausible MaSp repeat regions tailored to specific mechanical properties while also predicting those properties for given sequences. Validation includes sequence-level analysis, assessing physicochemical attributes and expected distribution of key motifs as well as secondary structure compositions. A correlation study using BLAST on the Spider Silkome dataset and a test set of MaSp repeats with known mechanical properties further confirmed the predictive accuracy of the model. This framework advances the rational design of spider silk-inspired biomaterials, offering a versatile tool for engineering protein sequences with tailored mechanical attributes.
@@ 1 Introduction
Recent advancements in protein design, particularly the integration of artificial intelligence (AI), have significantly enhanced our ability to engineer proteins with desired functions. Researchers have used deep learning techniques to improve the design of de novo proteins, achieving a tenfold increase in the success rates of target binding [1]. These innovations underscore the transformative potential of AI in protein engineering, paving the way for novel therapeutic interventions and biotechnological applications [2].
@@
Spiders spin up to seven different silk types, that all are composed of spidroins [3]. In this work, we focus on major ampullate silk, or dragline silk, renowned for exceptional mechanical properties - tensile strength comparable to steel (up to 1.3 GPa) with extensibility rivaling rubber (>30%) [4]. This unique combination yields toughness exceeding both steel and Kevlar [5], making spider silk particularly attractive for applications ranging from biomedical sutures to high-performance textiles [6].
@@
The molecular structure of MaSps comprises three primary regions: the globularly folded N-terminal and C-terminal domains, which are flanking the repetitive core region (Figure 1). The repetitive core region is believed to be the main contributor to the fiber mechanical properties, and in MaSps, this region is generally made up of poly-Ala blocks that are alternated with Gly-rich regions [7]. In the polymerized form, in the fiber, the poly-Ala blocks are predominantly arranged in nano-sized $\beta$-sheet crystals structures that contribute to the tensile strength of the silk fiber [8, 9, 10]. The poly-Ala $\beta$-sheet crystals are embedded in an amorphous matrix formed by the Gly-rich repeats, which is related to the extensibility of the fiber [11]. Although often referred to as amorphous, the Gly-rich repeats also form specific conformations like $3_{1}$-helices, $\beta$-turns and $\beta$-spirals [12, 13, 14].
@@
Native major ampullate silk fibers from different spider species display large variability in mechanical properties. To elucidate the source of this large variability, [15]. undertook the significant challenge of sequencing the transcriptome of 1098 species and simultaneously determining the mechanical properties of major ampullate silks from 446 of these species. The results showed that there is a large interspecies difference in terms of mechanical properties. For example, the tensile strength of the major ampullate silks varies between 0.17 and 3.3 GPa [15]. However, strong correlations between the amino acid sequence motifs in the repeat regions of the MaSps and the fiber mechanical properties could not be found (c.f. section 2.1).
@@
Efforts to produce artificial spider silk have largely focused on replicating natural silk spinning processes through a variety of engineered approaches [16, 17]. These methods majorly involve expression of engineered mini-spidroins in a heterologous host and subsequent spinning using an artificial spinning device [18]. Although these methods have significantly advanced synthetic production, they often fail to provide a scalable means to tailor silk properties for specific applications. This limitation further underscores the need for computational approaches to predict and design sequence–property relationships effectively.
In this study, we present a tailored multi-level strategy that addresses two key tasks: generating MaSp repeats customized for desired mechanical properties and predicting mechanical properties from given MaSp repeat sequences. Our model focuses exclusively on the repeat regions, distinguishing it from previous studies [19]. Our methodology begins with knowledge distillation of ProtGPT2 [20], a pre-trained generative model, by training a lightweight student model on 100k spider protein sequences, a subset of the UniRef50 database [21]. This distilled model then undergoes two-stage fine-tuning using the Spider Silkome dataset [15]: first on 6,000 MaSp repeat sequences to learn general repeat patterns, followed by refinement on 592 MaSp repeat sequences with known fiber-level mechanical properties.
@@
To evaluate the model’s effectiveness, we employed two distinct datasets: a test set of 20 instances sampled from the original 592-instance Spider Silkome dataset for assessing self-consistency, and a BLAST set curated to determine sequence novelty within a broader protein sequence database. Our evaluation followed a comprehensive two-level analysis approach. At the sequence level, we analyzed key properties including molecular weight, instability index, isoelectric point, and the distribution of essential motifs (GGX, poly-Ala, YGQGG, and SV). For structural validation, we employed a secondary structure prediction tool to verify the composition of $\alpha$-helices, $\beta$-strands and random coil characteristic of MaSps.
@@
  • A practical contribution to sustainable biomaterial development by providing a scalable approach for designing synthetic spider silk proteins with customizable mechanical properties.
@@ 2.1 Spider Silk: Sequence-property Relationships
Major ampullate silk is widely regarded a benchmark for high-performance biomaterials due to its unique combination of tensile strength, extensibility, and toughness [22, 5]. The fiber is primarily composed of MaSps, and to date, five distinct MaSp classes (MaSp 1-5) have been described [18]. The assignment of a MaSp to a specific class is determined by clustering of its terminal domains in phylogenetic analyses and the presence of characteristic amino motifs in the repetitive region.
Analyses of the spidroin sequences in the Spider Silkome database have identified several recurring amino acid motifs within the repetitive regions of the MaSps [15]. A number of these motifs have been highlighted in previous studies, including poly-Ala, GGX, YGQGG, SV, GPGXX, QQ, and AGQG [15, 23, 24, 25, 26, 27]. Attempts to link the occurrence of these motifs individually with the mechanical properties of major ampullate silks have only revealed weak or no correlations (Pearson correlation coefficients < 0.6) [15]. Noticeable among these are the YGQGG motif which is positively associated with toughness and the SV motif which is negatively correlated with this parameter [15]. The poly-Ala, GGX, QQ and AGQG motifs occur frequently in MaSps, but the occurrences of these motifs were not strongly correlated to the fiber mechanical properties [15]. This indicates that the sequence–function relationship of MaSps is complex and results from a synergistic interplay among multiple motifs.
In the present study, we leverage advancements in artificial intelligence to better understand the intricate sequence-to-property correlations, using data-driven models to unravel the hidden patterns governing silk mechanics.
@@ 2.2 Protein Design Using Machine Learning and Generative Models
Recent advances in artificial intelligence (AI) have significantly reshaped the landscape of protein sequence design, with generative pretrained language models (PLMs) playing a pivotal role. Models such as ProtGPT2 [20], ProGen [28], and ProtBERT [29] have harnessed natural language processing (NLP) techniques to generate and interpret biologically meaningful protein sequences. These models are trained on extensive datasets of protein sequences, enabling them to discern intricate patterns and relationships that underpin biological functionality [16]. Their success spans the design of enzymes, antibodies, and other functional proteins, frequently surpassing the capabilities of traditional directed evolution methods by producing sequences with enhanced activity or stability. Additionally, structure-aware models like AlphaFold2 [30] and ESMFold [31] have elevated sequence design by incorporating structural predictions, while emerging techniques, such as diffusion-based models [32], introduce innovative constraint-driven approaches for protein generation. In the context of spider silk proteins (spidroins), these generative PLMs are particularly promising due to their ability to capture the repetitive motifs and hierarchical organization inherent to spidroin sequences, though their application to optimizing mechanical properties remains an ongoing challenge.
@@ 2.3 Modeling Mechanical Properties of Spider Silk
Studies investigating the relationship between protein sequence and mechanical properties have primarily relied on sequence analysis and molecular simulations. For instance, [33] highlighted the importance of GGX and poly-Ala motifs in achieving spider silk’s extensibility and tensile properties. Secondary structure prediction tools like PSIPRED [34] have been widely used to predict $\alpha$-helices, $\beta$-strands and random coil arrangements,, which are critical for understanding silk mechanics [35]. However, these approaches are often retrospective and do not enable forward design of sequences with desired properties. Efforts to model mechanical properties using machine learning have shown promise. For instance, [36] utilized neural networks to predict silk’s mechanical properties based on amino acid composition. While such methods enhance understanding of sequence-property relationships, they lack the generative capability required for designing novel sequences. Recent advancements in hybrid AI techniques have shown promise in overcoming these limitations, particularly for engineering silk-like proteins with desired mechanical characteristics [19, 37, 38]. Notably, [19] introduced a generative modeling approach for spidroin sequence design, enabling the creation of synthetic spider silk proteins tailored to target mechanical properties. However, a key limitation of this approach is its lack of emphasis on MaSp repeat regions, which are the primary determinants of silk’s mechanical behavior.
@@ 2.4 Our Proposed Approach in Context
Existing studies have made significant strides in predicting and designing silk proteins, yet they often overlook the critical role of MaSp repeat regions in governing mechanical properties. While predictive models offer insights into sequence-property relationships, they lack generative capabilities, and current generative approaches fail to focus specifically on the functional repeats of MaSp. To address these gaps, our proposed approach introduces a multi-level fine-tuning strategy tailored for MaSp repeat generation. We first train a specialized model to learn the underlying structure and composition of MaSp repeats, ensuring the generated sequences remain biologically relevant. In the second stage, we fine-tune the model to establish a direct correlation between sequence patterns and mechanical properties, enabling controlled generation of MaSp repeats with desired mechanical characteristics. By integrating domain-specific fine-tuning with generative modeling, our method advances beyond existing approaches, providing a more precise and biologically grounded framework for spider silk protein design.
The ability to design spider silk proteins with tunable mechanical properties holds immense potential for sustainable material innovation. Applications extend from eco-friendly, biodegradable textiles to high-performance biomedical devices, including tissue scaffolds and targeted drug delivery systems [39]. Moreover, leveraging computational and experimental methodologies in protein engineering can accelerate the discovery of novel biomaterials, driving advancements in synthetic biology and biomimetic material design.
@@ 3 Proposed Pipeline: A Multi-Stage Fine-Tuning Framework
The target dataset, which maps MaSp repeats to their mechanical properties, is extremely small, posing a significant challenge for training generative deep neural networks. To overcome this limitation, we propose a multi-stage fine-tuning approach that enables the model to generate novel MaSp repeats for desired mechanical properties while also predicting mechanical properties from a given MaSp repeat. The training process is structured into three distinct stages, as illustrated in Figure 2. The following sections provide a detailed breakdown of each stage.
Figure 2: Illustration of the proposed methodology organized into three levels. Stage 1 involves training a distilled ProtGPT2 model using spider protein sequences from UniProtKB [21]. In Stage 2, the model is fine-tuned on the repeat regions of MaSp to adapt to their unique patterns. Finally, Stage 3 further fine-tunes the model to capture correlations between MaSp repeats and their mechanical properties.

### v2

@@ Abstract
The remarkable mechanical properties of spider silk, including its tensile strength and extensibility, are primarily governed by the repeat regions of the proteins that constitute the fiber, the major ampullate spidroins (MaSps). However, establishing correlations between mechanical characteristics and repeat sequences remains challenging due to the intricate sequence–structure–function relationships of MaSps and the limited availability of annotated datasets. In this study, we present a novel computational framework for designing MaSp repeat sequences with customizable mechanical properties. To achieve this, we developed a lightweight GPT-based generative model by distilling the pre-trained ProtGPT2 protein language model. The distilled model was subjected to multi-level fine-tuning using curated subsets of the Spider Silkome dataset. Specifically, we adapted the model for MaSp repeat generation using 6,000 MaSp repeat sequences and further refined it via cross-validation on 592 repeats associated with experimentally determined fiber-level mechanical properties. Our model generates biologically plausible MaSp repeat regions tailored to specific mechanical properties, while also predicting those properties for given sequences. Validation includes sequence-level analysis, assessing physicochemical attributes, the expected distribution of key motifs, and secondary structure compositions. A correlation study using BLAST on the Spider Silkome dataset and a test set of MaSp repeats with known mechanical properties further confirmed the predictive accuracy of the model. This framework advances the rational design of spider silk-inspired biomaterials, offering a versatile tool for engineering protein sequences with tailored mechanical attributes.
@@ 1 Introduction
Recent advancements in protein design, particularly the integration of artificial intelligence (AI), have significantly enhanced our ability to engineer proteins with desired functions. Researchers have used deep learning techniques to improve the design of de novo proteins, achieving a tenfold increase in the success rates of target binding (Bennett et al., 2023). These innovations underscore the transformative potential of AI in protein engineering, paving the way for novel therapeutic interventions and biotechnological applications (Khakzad et al., 2023).
@@
Spiders spin up to seven different silk types, that all are composed of spidroins (Peakall, 1969). In this work, we focus on major ampullate silk, or dragline silk, renowned for exceptional mechanical properties - tensile strength comparable to steel (up to 1.3 GPa) with extensibility rivaling rubber (>30%) (Vollrath & Knight, 2001). This unique combination yields toughness exceeding both steel and Kevlar (Gosline et al., 1999), making spider silk particularly attractive for applications ranging from biomedical sutures to high-performance textiles (Bourzac, 2015).
@@
The molecular structure of MaSps comprises three primary regions: the globularly folded N-terminal and C-terminal domains, which are flanking the repetitive core region (Figure 1). The repetitive core region is believed to be the main contributor to the fiber mechanical properties, and in MaSps, this region is generally made up of poly-Ala blocks that are alternated with Gly-rich regions (Guerette et al., 1996). In the polymerized form, in the fiber, the poly-Ala blocks are predominantly arranged in nano-sized $\beta$-sheet crystals structures that contribute to the tensile strength of the silk fiber (Hijirida et al., 1996; Lewis, 1992; Yarger et al., 2018b). The poly-Ala $\beta$-sheet crystals are embedded in an amorphous matrix formed by the Gly-rich repeats, which is related to the extensibility of the fiber (Gosline et al., 1984). Although often referred to as amorphous, the Gly-rich repeats also form specific conformations like $3_{1}$-helices, $\beta$-turns and $\beta$-spirals (Hayashi & Lewis, 1998; Hinman & Lewis, 1992; Van Beek et al., 2002).
@@
Native major ampullate silk fibers from different spider species display large variability in mechanical properties. To elucidate the source of this large variability, Arakawa et al. (2022). undertook the significant challenge of sequencing the transcriptome of 1098 species and simultaneously determining the mechanical properties of major ampullate silks from 446 of these species. The results showed that there is a large interspecies difference in terms of mechanical properties. For example, the tensile strength of the major ampullate silks varies between 0.17 and 3.3 GPa (Arakawa et al., 2022). However, strong correlations between the amino acid sequence motifs in the repeat regions of the MaSps and the fiber mechanical properties could not be found (c.f. section 2.1).
@@
Efforts to produce artificial spider silk have largely focused on replicating natural silk spinning processes through a variety of engineered approaches (Ferruz & Höcker, 2022; Koeppel & Holland, 2017). These methods majorly involve expression of engineered mini-spidroins in a heterologous host and subsequent spinning using an artificial spinning device (Schmuck et al., 2024). Although these methods have significantly advanced synthetic production, they often fail to provide a scalable means to tailor silk properties for specific applications. This limitation further underscores the need for computational approaches to predict and design sequence–property relationships effectively.
We propose a multi-stage framework for (1) generating MaSp repeat sequences conditioned on mechanical properties, and (2) predicting properties from given repeats. In contrast to prior work focused on full-length proteins (Lu et al., 2024), our model targets the repetitive core regions of MaSp proteins, which are central to fiber mechanics. The approach begins with distillation of ProtGPT2 (Ferruz et al., 2022) into a compact student model, SpiderGPT, trained on 100,000 spider proteins from UniRef50 (The UniProt Consortium, 2024). SpiderGPT is then fine-tuned in two stages using the Spider Silkome dataset (Arakawa et al., 2022): Level 1 uses 6,000 MaSp repeats to learn motif patterns, while Level 2 applies 5-fold cross-validation on 592 annotated sequences to learn sequence–property relationships. This hierarchical setup enables SpiderGPT to model both generative and predictive tasks with domain specificity and generalization.
@@
To evaluate the model’s effectiveness, we employed two distinct datasets: a test set of 185 instances sampled through cross-validation from the original 592-instance Spider Silkome dataset to assess self-consistency, and a BLAST set curated to determine sequence novelty against a broader protein sequence database. Our evaluation followed a comprehensive two-level analysis approach. At the sequence level, we analyzed key properties including molecular weight, instability index, isoelectric point, and the distribution of essential motifs (GGX, poly-Ala, YGQGG, and SV). For structural validation, we employed a secondary structure prediction tool to verify the composition of $\alpha$-helices, $\beta$-strands, and random coils characteristic of MaSps.
@@
  • A potential solution for the synthesis of sustainable biomaterials by providing a scalable approach for designing synthetic spider silk proteins with customizable mechanical properties.
@@ 2.1 Spider Silk: Sequence-property Relationships
Major ampullate silk is widely recognized as a benchmark for high-performance biomaterials due to its unique combination of tensile strength, extensibility, and toughness (Vollrath & Knight, 1999; Gosline et al., 1999). The fiber is primarily composed of major ampullate spidroins (MaSps), and to date, five distinct MaSp classes (MaSp1–MaSp5) have been identified (Schmuck et al., 2024). Assignment of a MaSp to a specific class is typically based on clustering of its terminal domains in phylogenetic analyses and the presence of characteristic amino acid motifs within the repeat region.
A typical MaSp sequence comprises three main components: an N-terminal domain, a repetitive core region, and a C-terminal domain, as illustrated in Figure 1. The terminal domains are highly conserved and shared between different spidroins that give rise to fibers of very diverse mechanical properties. This conservation suggests that terminal domains are primarily involved in spidroin storage and polymerization rather than in determining fiber mechanics. Both N- and C-terminal domains have been extensively characterized in prior work (Andersson et al., 2014; Askarieh et al., 2010; Eisoldt et al., 2012; Šede et al., 2022). To the best of our knowledge, direct interactions between terminal domains and the repeat region have not been reported. Moreover, evidence from the literature consistently supports the conclusion that the mechanical properties of spider silk fibers are governed predominantly by the repetitive core domains of the spidroins (Anton et al., 2017; Bowen et al., 2018; Keten & Buehler, 2010; Gosline et al., 1999; Kono et al., 2021a; Vollrath & Knight, 2001; Schmuck et al., 2024). Consequently, our design strategy focuses exclusively on the repeat region, targeting the sequence elements most directly responsible for material performance. While we do not exclude the possibility of terminal–repeat interactions, our methodological emphasis remains on the core motifs most relevant to mechanical function.
Analyses of spidroin sequences in the Spider Silkome database have revealed several recurring amino acid motifs within the repeat regions of MaSps (Arakawa et al., 2022). Notable motifs include poly-Ala, GGX, YGQGG, SV, GPGXX, QQ, and AGQG (Arakawa et al., 2022; Keten & Buehler, 2010; Malay et al., 2017; Craig et al., 2020; Nakamura et al., 2024b; Kono et al., 2021b). However, attempts to correlate individual motif occurrences with fiber mechanical properties have yielded weak or no relationships, with Pearson correlation coefficients generally below 0.6 (Arakawa et al., 2022). Among the few exceptions are the YGQGG motif, which shows a positive association with toughness, and the SV motif, which is negatively correlated with this parameter. Other frequently occurring motifs such as poly-Ala, GGX, QQ, and AGQG do not show strong individual correlations with fiber mechanics (Arakawa et al., 2022). These findings underscore the complexity of the sequence–function relationship in MaSps, which likely arises from synergistic interactions among multiple motifs rather than from the presence of any single motif.
In this study, we leverage recent advances in artificial intelligence to better understand the intricate sequence-to-property relationships within MaSp repeat regions. Through data-driven modeling, we aim to uncover latent patterns that govern the mechanical performance of spider silk fibers.
@@ 2.2 Protein Design Using Machine Learning and Generative Models
Deep learning has increasingly become a driving force in protein design, with generative models enabling data-driven exploration of protein sequence space. Among these, language model-based architectures have gained particular prominence by framing protein sequences analogously to natural language. Models such as ProtGPT2 (Ferruz et al., 2022), ProGen (Madani et al., 2020), and ProtBERT (Brandes et al., 2022) leverage transformer-based architectures trained on large-scale protein databases to generate biologically plausible sequences and predict structural or functional features.
These approaches have demonstrated success in designing enzymes (Madani et al., 2020), antimicrobial peptides (Das et al., 2021), and antibodies (Ruffolo et al., 2021), often surpassing traditional directed evolution methods in speed and diversity. More recently, models such as ProteinGAN (Repecka et al., 2021) and models conditioned on property constraints (Nijkamp et al., 2023; Lu et al., 2024) have emphasized functional design by guiding sequence generation toward specific biochemical or biophysical properties.
Structure-informed models like AlphaFold2 (Jumper et al., 2021), ESMFold (Lin et al., 2022), and diffusion-based frameworks (Watson et al., 2023; Anand et al., 2022) have further enhanced generative design by incorporating folding and energy priors, thus supporting the generation of not only functional but also structurally stable proteins.
In the context of spider silk proteins (spidroins), which are characterized by repetitive motifs and mechanical performance, these generative frameworks offer promising capabilities. However, sequence design aligned with fiber-level mechanical properties—such as toughness or strength—remains underexplored. Our work addresses this gap by conditioning sequence generation on mechanical property profiles, advancing property-aware generative modeling for hierarchical and repetitive proteins such as MaSp repeats.
@@ 2.3 Modeling Mechanical Properties of Spider Silk
Studies investigating the relationship between protein sequence and mechanical properties have primarily relied on sequence analysis and molecular simulations. For instance, (Tokareva et al., 2013) highlighted the importance of GGX and poly-Ala motifs in achieving spider silk’s extensibility and tensile properties. Secondary structure prediction tools like PSIPRED (McGuffin et al., 2000) have been widely used to predict $\alpha$-helices, $\beta$-strands and random coil arrangements,, which are critical for understanding silk mechanics (Buchan & Jones, 2019). However, these approaches are often retrospective and do not enable forward design of sequences with desired properties. Efforts to model mechanical properties using machine learning have shown promise. For instance, Kim et al. (2023) utilized neural networks to predict silk’s mechanical properties based on amino acid composition. While such methods enhance understanding of sequence-property relationships, they lack the generative capability required for designing novel sequences. Recent advancements in hybrid AI techniques have shown promise in overcoming these limitations, particularly for engineering silk-like proteins with desired mechanical characteristics (Lu et al., 2024; Ni et al., 2023; Fazio et al., 2023). Notably, Lu et al. (2024) introduced a generative modeling approach for spidroin sequence design, enabling the creation of synthetic spider silk proteins tailored to target mechanical properties. However, a key limitation of this approach is its lack of emphasis on MaSp repeat regions, which are the primary determinants of silk’s mechanical behavior.
@@ 2.4 Our Proposed Approach in Context
Existing studies have made significant progress in predicting and designing silk proteins; however, they often overlook the critical role of the repeat regions in MaSp that govern mechanical properties. In this study, we explicitly remove the terminal domains from MaSp sequences and focus our analysis solely on the repeat region.
Although predictive models have contributed to understanding sequence–property relationships, they typically lack generative capabilities. Conversely, existing generative approaches do not focus specifically on the functional repeats of MaSp. To bridge this gap, we introduce a multi-level fine-tuning strategy tailored for MaSp repeat generation. In the first stage, a specialized model is trained to learn the structural and compositional features of MaSp repeats, ensuring biological plausibility in generated sequences. In the second stage, we fine-tune the model on annotated sequences with mechanical property labels, establishing a direct correlation between sequence motifs and their functional outcomes. This enables controlled generation of MaSp repeats with tunable mechanical characteristics.
By integrating domain-specific fine-tuning with generative modeling, our method extends beyond existing approaches, providing a more precise and biologically grounded framework for spider silk protein design. The ability to design spider silk proteins with tunable mechanical properties has significant implications for sustainable material innovation. Applications include eco-friendly, biodegradable textiles and high-performance biomedical materials such as tissue scaffolds and drug delivery systems (De Giorgio et al., 2024). More broadly, the combination of computational and experimental tools in protein engineering promises to accelerate the discovery of novel biomaterials and advance the field of biomimetic material design.
@@ 3 Proposed Pipeline: A Multi-Stage Fine-Tuning Framework
The target dataset, which maps MaSp repeats to their mechanical properties, is extremely limited in size, presenting a major challenge for training generative deep neural networks. To address this limitation, we adopt a multi-stage fine-tuning strategy that enables the model to perform two complementary tasks: (1) generating novel MaSp repeat sequences conditioned on desired mechanical properties, and (2) predicting mechanical properties from a given MaSp repeat.
The training pipeline is organized into three stages, as outlined in Figure 2, and is designed to progressively distill and specialize domain knowledge across tasks and data scales. While this hierarchical strategy is tailored to protein design, similar multi-stage or two-step post-training approaches have been explored in other domains. For instance, a recent study in audio modeling applies a conceptually similar two-phase training scheme for waveform generation (Sinha et al., 2024), described as a “two-step post-training” process.
The following sections detail the implementation and rationale of each stage within our framework.
Figure 2: Illustration of the proposed methodology organized into three levels. Stage 1 involves training a distilled ProtGPT2 model using spider protein sequences from UniProtKB (The UniProt Consortium, 2024). In Stage 2, the model is fine-tuned on the repeat regions of MaSp to adapt to their unique patterns. Finally, Stage 3 further fine-tunes the model to capture correlations between MaSp repeats and their mechanical properties.